OpenStudy (anonymous):

what is the answer of log base 5 of root 10..?? plz rply

OpenStudy (anonymous):

\[\log _{5} \sqrt{10} ?\]

OpenStudy (ash2326):

@sana_kausar Can you simplify this?

OpenStudy (anonymous):

@mrbarry... yes what's d answer fr it..??

ganeshie8 (ganeshie8):

it has to be less than 1 cuz \(\sqrt{10}\) is less than 5

OpenStudy (unklerhaukus):

\[\log_5 \sqrt{10}=\log_5 10^{1/2}=\tfrac12\log_5 10=\tfrac12\log_5 (5\times2)=\]

OpenStudy (anonymous):

\[\sqrt{10} is 0.5\]

OpenStudy (primeralph):